solve the following homogeneous system of linear equations:
x+y-3y+2w=0,
2x-y+2z-3w=0,
3x-2y+z-4w=0,
-4x+y-3z+w=0

Answer by Alan3354(69443)

solve the following homogeneous system of linear equations:
x+y-3y+2w=0,
2x-y+2z-3w=0,
3x-2y+z-4w=0,
-4x+y-3z+w=0
--------------
All rows = 0
---
w = x = y = z = 0